OpenStudy (anonymous):

What is the slope of a line parallel to the line whose equation is 5x + 3y = 15?

OpenStudy (anonymous):

slope of \[ax+by=c\] is \[\frac{-a}{b}\]

OpenStudy (anonymous):

@Nnesha Can you help me understand how to solve this

OpenStudy (anonymous):

I plugged in what he gave me and i got -1.666666666666667

Nnesha (nnesha):

ok so you have to solve for y

Nnesha (nnesha):

5x + 3y = 15 what is your first step suppose to be ???

OpenStudy (anonymous):

Ohh. Wait I think I remember this. I would replace x with 0 right? So 5*0 + 3y = 15 so 0 + 3y = 15, so then I divide by 3? y = 5?

Nnesha (nnesha):

nope you don't have to plug 0 you have to keep the x we need equation in slope intercept form like y = mx + b

OpenStudy (anonymous):

Ohh. Then I don't know what the first step would be

Nnesha (nnesha):

you have to move 5x on right side 5x is adding on left side what is opposite of add ??

OpenStudy (anonymous):

subtract

Nnesha (nnesha):

yes so subtract 5x both side

OpenStudy (anonymous):

5x + 3y = 15 -5x -5x 3y = 20x

Nnesha (nnesha):

you cannot combine 15 and -5x because -5 have variable and 15 doesn't so leave it as 3y = -5x + 15

Nnesha (nnesha):

now divide both side by 3

OpenStudy (anonymous):

y = -5x + 5?

Nnesha (nnesha):

-5x / 3 = ???

OpenStudy (anonymous):

well it equals -1.666666666666667 according to my calculater

Nnesha (nnesha):

leave it as fraction form so it should be -5x/3 not just -5x so your equation is y = -5x/3 + 5

OpenStudy (anonymous):

Oh ok. So then what

OpenStudy (anonymous):

I have to leave. we can continue this tomorrow when you're available :)

Nnesha (nnesha):

so parallel lines have same slope

Nnesha (nnesha):

we can solve now we almost done :)

OpenStudy (anonymous):

im ready when you are :)

Nnesha (nnesha):

yep so parallel lines have same slope

Nnesha (nnesha):

what is slope of that equation

OpenStudy (anonymous):

-5?

Nnesha (nnesha):

y = -5x/3 + 5 your slope is -5/3 so because other line is parallel then slope gonna be same which is -5/3

OpenStudy (anonymous):

YES! !Thank you